Earth (mass = 6.00 × 1027 g) moving through space at 3.20 × 104 m/s. Answer in nm.
1
Expert's answer
2019-10-01T03:24:48-0400

As momentum "(p)" "=\frac{h(Js)}{wavelength(m)}"

wavelength"=\frac{h}{p}" "=\frac{h}{mv}" "=\frac{6.626\times10^{-34}}{6\times10^{24}Kg\times3.20\times10^4}=0.345\times10^{-62}m"

Hence, wavelength in "nm=0.345\times10^{-62}\times10^9=0.345\times10^{-53}nm"